OpenStudy (anonymous):

Hi everyone! Using u=y-2x+3, Solve dy/dx=2+(root(y-2x+3))...Anyone? Kinda lost here :o)

OpenStudy (anonymous):

I get du/dx = -2dy/dx and -1/2du/dx=dy/dx...I have tried a bunch of stuff but can't figure it out.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

u=y-2x+3 y is a function of x here, so the derivative of y with respect to x will not be 0

ganeshie8 (ganeshie8):

\[u = y-2x+3\] \[\implies \frac{du}{dx} = \frac{dy}{dx} - 2\]

OpenStudy (anonymous):

how do you do that? I thought you take the partial multiplied by dy/dx?

OpenStudy (anonymous):

I got -2dy/dx

OpenStudy (anonymous):

what am I missing here?

OpenStudy (ribhu):

@SinginDaCalc2Blues i have tried to solve it for you, if you could get it.

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

okay, let me take a look! :O)

OpenStudy (ribhu):

sorry i messed it in the last step @SinginDaCalc2Blues

OpenStudy (ribhu):

I messed the integration, the final answer would be \[2\sqrt{y-2x+3}\] = x+c

OpenStudy (ribhu):

@SinginDaCalc2Blues just see it

OpenStudy (anonymous):

It looks like it may be wrong...I'm pretty sure the integral of 1/root(u) = 2root(u), so I think the final solution should be x - 2root(y-2x+3) + c =0 Can you check your work to make sure?

OpenStudy (ribhu):

yeah thats what i wrote after correcting myself and i did admit that i messed up the integration.

OpenStudy (anonymous):

Oh, I didn't see your post. Thank you very much for the help! :o)

OpenStudy (ribhu):

u got my handwriting clearly?? lol

OpenStudy (anonymous):

yes thanks